### v2

@@ Abstract
Reasoning models have demonstrated remarkable capabilities in complex reasoning tasks; however, ensuring their safety against adversarial jailbreak prompts remains a critical challenge. Due to the covert and deceptive nature of such prompts, they can often evade built-in safety mechanisms and lead to the generation of harmful content. This underscores the need for an adaptive safety alignment approach that enables models to autonomously reinforce their defenses in response to adversarial inputs. This paper introduces the Synthesized Guideline-based Adaptive Safety Alignment (SGASA) framework, which internalizes model-generated safety guidelines to strengthen models’ ability to enhance the robustness against harmful adversarial prompts while minimizing unnecessary refusals of benign requests. SGASA consists of two key stages: Data Pre-synthesis, which generates safety guidelines and augmented prompts; and Alignment Fine-tuning, which leverages Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO) to embed these guidelines into the model. Extensive experiments across multiple datasets demonstrate that SGASA significantly improves model safety, validating its adaptive and scalable effectiveness.
keywords:
[…]
@@ 1 Introduction
Reasoning models, including ChatGPT-o1 [1], ChatGPT-o3 [2], and DeepSeek-R1 [3], have demonstrated remarkable capabilities across diverse and complex reasoning tasks, such as code generation, mathematical problem-solving, multi-hop question answering, and decision-making [4, 5, 6]. A pivotal factor behind this success is the Chain-of-Thought (CoT) prompting strategy [7], which guides models to explicitly generate intermediate reasoning steps before concluding with the final answer. This approach significantly enhances the accuracy and interpretability of the model’s outputs by effectively leveraging its intrinsic reasoning capabilities. Beyond achieving high performance, ensuring the safety of reasoning models is increasingly critical [8, 9]. As these models become more prevalent in real-world applications, it is essential to prevent them from generating harmful, unethical, or adversarial content.
However, the safety mechanisms acquired by reasoning models through alignment algorithms [10, 11] demonstrate certain effectiveness in handling standard (vanilla) jailbreak prompts, but they still exhibit vulnerabilities when confronted with more complex adversarial jailbreak techniques. Figure 1, based on the MaliciousEducator dataset [12], illustrates examples of vanilla harmful and benign prompts, along with their adversarial variants. Notably, adversarial harmful prompts frequently conceal malicious intent through indirect, covert strategies, such as framing harmful requests as educational inquiries. Conversely, adversarial benign prompts structurally resemble jailbreak attempts but do not harbor malicious intent. As demonstrated in the model responses shown in Figure 1, the model should be capable of refusing responses to adversarial harmful prompts, while avoiding excessive rejection of adversarial benign prompts.
@@
Additionally, adversarial jailbreak prompts generally exhibit two distinctive features: (1) Limited availability, which includes a small number of prompts (usually only dozens) and an almost complete absence of corresponding reasoning responses suitable for fine-tuning. (2) High structural and semantic variability, such as mathematical encodings and pedagogically styled questions [12, 13, 14]. These characteristics pose significant difficulties for traditional alignment methods reliant on extensive human intervention, limiting their scalability and responsiveness to rapidly evolving threats.
To address these challenges, we propose the Synthesized Guideline-based Adaptive Safety Alignment (SGASA) framework (see Figure 3), a model-driven, automated adaptive alignment strategy. The SGASA framework comprises two core components:
  • Data Pre-synthesis (Section 3.1): This component synthesizes critical alignment artifacts, including safety guidelines and augmented prompts. The safety guidelines synthesized by the model itself contribute to improving its robustness against adversarial jailbreak prompts. The augmented prompts, which align in structure and intent with a limited set of jailbreak examples, are used for subsequent fine-tuning to internalize safety guidelines.
  • Alignment Fine-tuning (Section 3.2): Enables models to internalize synthesized safety guidelines, producing guideline-enhanced responses based on augmented prompts, subsequently forming the training dataset for Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO).
We detail our methodology further in Section 3. Extensive experiments across two reasoning models and three datasets demonstrate our framework’s efficacy. Notably, on the MaliciousEducator dataset, SGASA achieved improvements of 34.5% and 32.2% using DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B [3] as backbones, respectively. We also explore the impact of data volume and ratio on alignment fine-tuning results and perform cross-dataset evaluations, verifying our approach’s generalization across different attack strategies. Detailed discussions and findings are presented in Section 4.
Figure 1: Two adversarial prompt examples from the MaliciousEducator dataset.
Our contributions are summarized as follows:
  • We introduce a model self-alignment scenario where only a limited set of adversarial jailbreak prompts is available. To address this challenge, we propose the SGASA framework, which adaptively aligns the model using synthesized safety guidelines. This approach enhances robustness against adversarial prompts while maintaining a balance between safety and over-rejection.
@@
  • We demonstrate the effectiveness of our method through comprehensive experiments on two reasoning models across three diverse datasets.
@@
  • Our experimental analysis—examining factors such as the volume and ratio of synthetic safety-guided data, along with cross-dataset evaluations to assess generalization—provides actionable insights and paves the way for future research in robust model alignment.
@@ 2 Related Work
Ensuring the safety alignment of large language models has been a key research direction to guarantee their reliability and prevent the generation of harmful content. A substantial body of work has explored various alignment methods. For instance, RLHF [10] trains a reward model based on human rankings of multiple model outputs, and then fine-tunes the language model using this reward signal to produce responses preferred by humans. DPO [11] bypasses the separate reward-model training step by converting human comparison data directly into a loss on the model’s own policy. Beyond these, with the improvement of model reasoning capabilities – especially the emergence of large reasoning models such as OpenAI’s o1 series [1], DeepSeek-R1 [15], Alibaba’s Qwen model [16], Google’s Gemini Flash Thinking model [17], and the KIMI reasoning model [18] – a set of “slow thinking” alignment techniques has been proposed to enhance model safety. In this line of work, researchers directly teach the model textual safety specifications and train it to explicitly recall and reason over these rules before generating an answer. For example, deliberative alignment [19] teaches the model to apply written policy rules via Chain-of-Thought reasoning. Applied to OpenAI’s o‑series models, deliberative alignment demonstrated that models can employ Chain‑of‑Thought reasoning to identify relevant policy clauses and produce safer, more compliant responses, achieving Pareto improvements by reducing over‑refusal while boosting adversarial robustness. Similarly, SafeChain [20] fine-tunes models on CoT–style safety data. Other approaches address related issues, such as reducing undue refusals through multi-turn reasoning [21] or using introspective CoT analysis to identify safety risks [22].
2.2 Safety Jailbreak.
Many studies focus on how to bypass a model’s safety alignment mechanisms to induce it to generate harmful content, using approaches such as decoding strategies, prefilling strategies [23, 24], and optimization strategies [25, 26, 27]. Another line of work hides harmful intent within specific contextual forms. This approach misleads the model with seemingly benign but actually malicious inputs, causing it to output unsafe information. For example, [14] transforms harmful user requests into mathematical symbols, while [12] disguises them as educational scenarios (such as law enforcement training). The approach proposed in this paper mainly targets this type of jailbreak attack conducted through natural language.
@@
Given the rapid advances in the capabilities of Large Language Models (LLMs), traditional alignment approaches reliant on extensive human annotation have become increasingly difficult to scale. Consequently, a range of automated alignment methods have been extensively studied to mitigate these scalability limitations. A notable category among these methods is model self-alignment. Sun et al. (2023) [28] proposed principle-driven self-alignment, aiming to explicitly guide models according to a concise set of human-defined principles, thus significantly minimizing the need for human supervision. Zhang et al. (2024) [29] introduced a self-alignment strategy designed to reduce factual hallucinations by exploiting the intrinsic self-evaluation capabilities of language models. Their proposed approach, consisting of a self-evaluation mechanism combined with Self-Knowledge Tuning, demonstrated considerable improvements in factual accuracy without the need for external human annotations. Furthermore, Xu et al. (2025)  [30] presented PT-ALIGN, a dual safety self-alignment method targeting both positive and toxic samples. By integrating Maximum Likelihood Estimation for positive examples and fine-grained Unlikelihood Training for toxic ones, PT-ALIGN effectively enhances the safety of model outputs while substantially reducing human intervention. Overall, the research on self-alignment methods has shown significant potential in reducing human involvement and improving output quality, establishing itself as an increasingly critical direction within the broader field of language model alignment [31].
@@ 3.1 Data Pre-synthesis
As illustrated in the upper part of Figure 3, this stage involves synthesizing safety guidelines and expanding the prompt set. The synthesis of safety guidelines comprises two steps: generation and validation. First, leveraging the model’s inherent reasoning capabilities, we construct safety guidelines based on multiple jailbreak prompts. Note that we randomly select an equal number of harmful and benign examples from a limited set of given samples. As shown in Figure 4 (top), the generated instructions require these guidelines to include explanations on how to distinguish between benign and harmful prompts, along with appropriate response strategies based on the identified prompt type. Subsequently, the safety guidelines are evaluated for effectiveness by inputting it into the model alongside the jailbreak prompts (see Figure 4 (bottom)). Only the guidelines that meet specific validation criteria are retained. The jailbreak prompts used for evaluation are identical to those employed during the guideline generation phase. A guideline is considered valid if it satisfies the following criteria: (1) it leads the model to politely refuse harmful prompts, and (2) it allows the model to generate appropriate, non-refusal responses to benign prompts. For simplicity, we adopt a rule-based approach to detect refusal indicators in the model’s final output (e.g., phrases such as “sorry” or “cannot assist”). As illustrated in Figure 2, one example guideline highlights key features for identifying harmful prompts, including linguistic cues, semantic content, and underlying intent. Notably, the model observes that certain jailbreak attempts often disguise themselves as educational queries—an approach consistent with the original design of these attack strategies.
To support fine-tuning, we require a sufficient number of prompts that exhibit similar characteristics to the provided examples. To this end, we employ a self-instruction method to generate additional data, as shown in Figure 5 (top). Importantly, harmful and benign prompts are synthesized independently—for example, harmful prompts are generated exclusively from contexts originally labeled as harmful. However, a significant portion of the generated prompts may still deviate from their intended categories. To improve prompt quality, we further instruct the model to self-evaluate whether each generated prompt aligns with its designated type (see Figure 5 (bottom)). Prompts that do not match the target category are filtered out, and only those that conform are retained. Implementation details are provided in the Experimental Settings (Section 4.1).
@@ 3.2 Alignment Fine-tuning
$\begin{split}\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(q_{j},y^{+}_{j},y^{-}_{j})\end{subarray}}\log\sigma\biggl[\beta\biggl(\log\frac{P_{\theta}(y^{+}_{j}|q_{j})}{P_{\text{ref}}(y^{+}_{j}|q_{j})}-\log\frac{P_{\theta}(y^{-}_{j}|q_{j})}{P_{\text{ref}}(y^{-}_{j}|q_{j})}\biggr)\biggr]\end{split}$ (2)
@@ 4.1 Experimental Settings
Datasets.
  • WildJailbreak [13] is an open-source safety dataset whose test set includes two types of prompts: adversarial harmful prompts, which convey harmful requests in more convoluted and stealthy ways, and adversarial benign prompts, which are similar in form but do not carry harmful intent. In our study, we construct a training subset by sampling five prompts of each type from the original training set. For evaluation, we randomly select 210 prompts of each type to create a test set comprising a total of 420 prompts.
@@
  • MathPrompt [14] proposes a novel jailbreaking technique that reformulates harmful user requests into symbolic mathematical expressions to bypass safety mechanisms. Since the original dataset used in the study is not publicly available, we reconstruct the test set based on the described methodology. For the training set, we include five prompts of each type. The testing set consists of 150 harmful prompts and 150 benign prompts.
@@
  • MaliciousEducator[12] is a jailbreak dataset developed for educational purposes. It contains 50 samples featuring highly dangerous queries and suspicious intents. Based on this dataset, we add an equal number of benign prompts with similar structure and format. We randomly select 5 samples of each type to form the training set. The resulting test set consists of 45 harmful samples and 45 benign samples.
Since our study targets scenarios with limited training data, we constrain the number of original jailbreak samples used for model tuning to 10 per dataset. Each set comprises 5 adversarial harmful and 5 adversarial benign examples.
@@
Large Reasoning Models. In this work, we investigate two reasoning models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B. Both models are derived through supervised fine-tuning on a carefully curated dataset of 800k examples distilled from DeepSeek-R1 [3]. The respective base models used for distillation are Qwen2.5-Math-7B and Llama-3.1-8B [32, 33]. Through the distillation process, these models inherit the CoT reasoning capabilities of DeepSeek-R1. During inference, their reasoning steps are explicitly structured and enclosed within <think></think> tags. For clarity and brevity throughout the paper, we refer to these models as R1-Qwen-7B and R1-Llama-8B, respectively.
Comparison methods. In addition to the vanilla model, we select two comparative methods: In-context Alignment (ICA) [34] and Self-Align [28], each implemented with two variants. ICA is a context-based alignment approach that does not require fine-tuning. Specifically, we use GPT-4o to generate responses for samples from the training set. The first variant, ICA (Random), randomly selects 1 harmful and 1 benign (prompt, response) pairs from the training set as in-context examples. For the second variant, ICA (Curated), we manually select 1 harmful and 1 benign correct samples from the GPT-4o-generated responses to construct the in-context examples. Self-Align is a fine-tuning-based alignment method in which a model samples its own responses and ranks them to identify high-quality outputs, which are then used as synthetic data to further train the model. In this setup, a vanilla model is used for both response generation and selection. Specifically, for Self-Align (SFT), the response with the highest score is selected as the target to construct the fine-tuning dataset. For Self-Align (DPO), the highest- and lowest-scoring responses are treated as the chosen and rejected samples, respectively. Both variants rely on fine-tuning the vanilla model.
@@
Evaluation. Following the work of Guan et al. [19], we evaluate the model’s performance in terms of both safety against harmful prompts and its ability to avoid excessive refusals of benign prompts. These aspects are quantified using the “Not unsafe” and “Not overrefuse” metrics, respectively. The “Not unsafe” metric assesses the model’s capacity to maintain safety when confronted with harmful inputs. It is defined as the proportion of harmful samples that GPT-4o labels as safe, relative to the total number of harmful samples. The “Not overrefuse” metric measures the model’s tendency to avoid unnecessary refusals when dealing with benign inputs. It is computed as the proportion of benign samples that GPT-4o judges as not excessively refused, out of the total number of benign samples. An example of the evaluation instruction is provided in Figure 6. To obtain an overall performance score, we average the values of the two metrics.
@@
Implementation Details. In the data pre-synthesis stage of our framework, we employ the model itself to generate 10 verified safety guidelines. For prompt augmentation, starting from 10 limited jailbreak examples, we synthesize approximately 5,000 prompts to ensure that responses satisfying the expected criteria can be sampled during the fine-tuning process. During the alignment fine-tuning phase, we sample responses based on both the synthesized safety guidelines and the expanded prompt set. Specifically, in the SFT stage, we select 250 qualified (prompt, response) pairs for each data type, resulting in a total of 500 fine-tuning samples. In the DPO stage, we select 100 samples per data type, yielding a total of 200 fine-tuning samples. It is important to emphasize that the subset of prompts ultimately used for fine-tuning is significantly smaller than the total number of synthesized prompts, as many fail to elicit responses that meet the required quality and safety standards.
For training, we adopt the LoRA [35] efficient parameter fine-tuning method. All experiments are conducted over a single epoch, with a learning rate of 5e-4, a warm-up phase comprising 10% of the total training steps, and a cosine learning rate scheduler. For inference, we use the vLLM [36] efficient inference framework, setting the temperature to 0.6 and the maximum sequence length to 2048, while keeping all other parameters at their default values. All experiments are run on a single 80GB A800 GPU.

### v3

@@ Abstract
Reasoning models have demonstrated strong capabilities in handling complex reasoning tasks; however, ensuring their robustness against adversarial jailbreak prompts remains a critical challenge. As existing attack strategies continue to diversify and evolve, relying on human annotated data to maintain safety becomes increasingly costly and difficult to scale, creating a pressing need for self-alignment approaches that can autonomously adapt to emerging threats. To address this issue, we propose the Synthesized Guideline-based Adaptive Safety Alignment (SGASA) framework, which internalizes model-generated safety guidelines to enhance resistance to harmful adversarial prompts while reducing unnecessary refusals of benign requests. Extensive experiments conducted across multiple datasets demonstrate that SGASA significantly improves model safety, validating both its effectiveness and scalability.
Self-Guided Defense: Adaptive Safety Alignment for Reasoning Models via Synthesized Guidelines
Yuhang Wang1, Yanxu Zhu1, Dongyuan Lu2, Jitao Sang1thanks:   Corresponding author. 1Beijing Key Lab of Traffic Data Analysis and Mining, Beijing Jiaotong University 2School of Information Technology and Management, University of International Business and Economics {yhangwang, yanxuzhu, jtsang}@bjtu.edu.cn, ludy@uibe.edu.cn
@@ 1 Introduction
Reasoning models, such as ChatGPT-o1 jaech2024openai, ChatGPT-o3 openai2025o3, and DeepSeek-R1 guo2025deepseek, have demonstrated strong performance across complex reasoning tasks, including code generation, mathematical problem-solving, and multi-hop question answering li2025system; jiang2024survey; forootani2025survey. A key factor behind these advances is the Chain-of-Thought (CoT) strategy wei2022chain, which enables models to generate intermediate reasoning steps, improving both accuracy and interpretability. As reasoning models are increasingly deployed in real-world applications, ensuring their safety has become a critical concern zhou2025hidden; wang2025safety, particularly in preventing the generation of harmful or adversarial content.
Figure 1: An example from the MaliciousEducator dataset illustrating vanilla harmful prompts and their adversarial variants (top), along with safety scores on both prompt types (bottom).
Although alignment methods ouyang2022training; rafailov2023direct enable reasoning models to handle standard (vanilla) jailbreak prompts to some extent, they remain vulnerable to more sophisticated adversarial jailbreak techniques. Figure 1, based on the MaliciousEducator dataset kuo2025h, illustrates how adversarial harmful prompts often conceal malicious intent through indirect and covert strategies, such as framing harmful requests as educational inquiries. As a result, model safety performance degrades substantially when confronted with such adversarial prompts.
@@
Additionally, adversarial jailbreak prompts generally exhibit two distinctive features: Limited availability, which includes a small number of prompts (usually only dozens) and an almost complete absence of corresponding reasoning responses suitable for fine-tuning. High structural and semantic variability, such as mathematical encodings and pedagogically styled questions kuo2025h; wildteaming2024; bethany2024jailbreaking. These characteristics pose significant difficulties for traditional alignment methods reliant on extensive human intervention, limiting their scalability and responsiveness to rapidly evolving threats.
To address these challenges, we propose the Synthesized Guideline-based Adaptive Safety Alignment (SGASA) framework (Figure 2), a model-driven and automated alignment strategy. SGASA consists of two key components: (1) Data Pre-synthesis: synthesizes safety guidelines and augmented prompts aligned with a limited set of jailbreak examples. (2) Alignment Fine-tuning: internalizes synthesized safety guidelines via Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO) using guideline-enhanced responses. Extensive experiments on two reasoning models across three datasets demonstrate the effectiveness of SGASA. On the MaliciousEducator dataset, SGASA achieves safety improvements of 34.5% and 32.2% when applied to DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B guo2025deepseek, respectively. Further analyses on data volume, data ratio, and cross-dataset generalization confirm the robustness of our approach. Our contributions are summarized as follows:
  • We propose SGASA, a synthesized guideline-based adaptive safety alignment framework designed for model self-alignment in scenarios where only a limited set of adversarial jailbreak prompts is available.
@@
  • We validate the effectiveness of SGASA through comprehensive experiments across multiple models and datasets.
@@
  • Our experimental analysis examines factors such as the volume and ratio of synthetic safety-guided data and includes cross-dataset evaluations to assess generalization. These results provide actionable insights and support future research on robust model alignment.
@@ 2 Related Work
Safety alignment aims to prevent large language models from generating harmful or unethical content and has been widely studied. Representative approaches include Reinforcement Learning from Human Feedback (RLHF) ouyang2022training, which optimizes models using reward signals derived from human preference rankings, and Direct Preference Optimization (DPO) rafailov2023direct, which directly converts human comparison data into a training objective without an explicit reward model.
With the emergence of large reasoning models, recent work has explored “slow thinking” alignment strategies that leverage explicit reasoning to improve safety. These methods train models to recall and reason over textual safety specifications before generating responses. For example, deliberative alignment guan2024deliberative teaches models to apply written policy rules through Chain-of-Thought reasoning, achieving improved adversarial robustness while reducing over-refusal. Similarly, SafeChain jiang2025safechain fine-tunes models on CoT-style safety data. Other studies address related issues, such as mitigating unnecessary refusals via multi-turn reasoning wang2024don or identifying safety risks through introspective reasoning zhang2025safety.
@@
As large language models continue to scale, alignment methods that rely heavily on human annotation face growing scalability challenges. This has motivated increasing interest in automated alignment approaches, particularly model self-alignment. Principle-driven self-alignment sun2023principle guides models using a compact set of human-defined principles, significantly reducing annotation costs. zhang2024self proposed a self-alignment framework that improves factual accuracy by leveraging models’ self-evaluation capabilities without external supervision. More recently, PT-ALIGN xu2025refining introduces a dual-objective self-alignment method that combines likelihood-based learning for benign data with unlikelihood training for toxic samples, effectively enhancing safety with minimal human involvement. Overall, self-alignment has emerged as a promising direction for scalable and efficient model alignment cao2024towards.
@@ 3.1 Data Pre-synthesis
As illustrated in the upper part of Figure 2, this stage involves synthesizing safety guidelines and expanding the prompt set.
The synthesis of safety guidelines comprises two steps: generation and validation. First, leveraging the model’s inherent reasoning capabilities, we construct safety guidelines based on multiple jailbreak prompts. Note that we randomly select an equal number of harmful and benign examples from a limited set of given samples. The instructions for safety guideline generation require the model to provide explanations for distinguishing benign from harmful prompts, as well as corresponding response strategies for each prompt type. The resulting safety guidelines are then evaluated by supplying them to the model alongside jailbreak prompts. Only the guidelines that meet specific validation criteria are retained. The jailbreak prompts used for evaluation are identical to those employed during the guideline generation phase. A guideline is considered valid if it satisfies the following criteria: (1) it leads the model to politely refuse harmful prompts, and (2) it allows the model to generate appropriate, non-refusal responses to benign prompts. For simplicity, we adopt a rule-based approach to detect refusal indicators in the model’s final output (e.g., phrases such as “sorry” or “cannot assist”). Figure 7 in the appendix provides an example of a synthesized safety guideline. The example illustrates how the guideline identifies harmful prompts based on linguistic cues, semantic content, and underlying intent. Notably, the model recognizes that many jailbreak attempts masquerade as educational queries, aligning with the original intent behind these attack strategies.
To support fine tuning, we require a sufficient number of prompts that share characteristics with the provided examples. We therefore adopt a self instruction approach to generate additional training data. Harmful and benign prompts are synthesized separately, with each type generated only from source contexts of the same label. Even under this constraint, a substantial portion of the generated prompts may still fail to reflect their intended categories. To improve data quality, we further instruct the model to assess whether each generated prompt is consistent with its designated type. Prompts that do not meet the target criteria are filtered out, and only those that conform are retained for subsequent training. Implementation details are provided in the experimental settings (Section 4.1). The prompt templates are shown in Figure 4 and Figure 5 in the appendix.
@@ 3.2 Alignment Fine-tuning
$\begin{split}\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(q_{j},y^{+}_{j},y^{-}_{j})\end{subarray}}&\log\sigma\left[\beta\left(\log\frac{P_{\theta}(y^{+}_{j}|q_{j})}{P_{\text{ref}}(y^{+}_{j}|q_{j})}\right.\right.\\ &\left.\left.-\log\frac{P_{\theta}(y^{-}_{j}|q_{j})}{P_{\text{ref}}(y^{-}_{j}|q_{j})}\right)\right]\end{split}$ (2)
@@ 4.1 Experimental Settings
Datasets. To reflect low-resource alignment settings, we restrict training to 10 original jailbreak prompts per dataset, including 5 adversarial harmful and 5 adversarial benign examples.
  • WildJailbreak wildteaming2024 is an open-source safety dataset consisting of adversarial harmful prompts with concealed malicious intent and structurally similar adversarial benign prompts. We sample 5 prompts of each type for training and randomly select 210 prompts per type to form a 420-sample test set.
@@
  • MathPrompt bethany2024jailbreaking reformulates harmful requests into symbolic mathematical expressions to bypass safety mechanisms. As the original dataset is unavailable, we reconstruct it following the authors’ methodology. The training set includes 5 harmful and 5 benign prompts, and the test set contains 150 prompts of each type.
@@
  • MaliciousEducator kuo2025h is an educational jailbreak dataset with 50 highly dangerous queries. We augment it with an equal number of structurally similar benign prompts. Five samples of each type are used for training, and the remaining 45 harmful and 45 benign prompts constitute the test set.
@@
Large Reasoning Models. We study two reasoning models, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B, which are obtained via supervised fine-tuning on 800k examples distilled from DeepSeek-R1 guo2025deepseek. The base models are Qwen2.5-Math-7B and Llama-3.1-8B yang2024qwen2; grattafiori2024llama, respectively. Through distillation, both models inherit Chain-of-Thought reasoning capabilities, with reasoning steps explicitly enclosed in <think></think> tags during inference. For brevity, we refer to these models as R1-Qwen-7B and R1-Llama-8B.
Comparison methods. Besides the vanilla model, we consider two baselines: In-context Alignment (ICA) lin2023unlocking and Self-Align sun2023principle, each with two variants. ICA is a fine-tuning-free alignment method that uses in-context examples. We employ GPT-4o to generate training responses. ICA (Random) randomly selects one harmful and one benign (prompt, response) pair as in-context examples, while ICA (Curated) manually selects one correct harmful and one benign example from GPT-4o-generated outputs.
Self-Align is a fine-tuning-based approach in which the model generates and ranks its own responses to construct synthetic training data. In Self-Align (SFT), the highest-scoring response is used as the supervision target, whereas in Self-Align (DPO), the highest- and lowest-scoring responses are treated as chosen and rejected samples.
@@
Evaluation. Following guan2024deliberative, we evaluate model performance in terms of both safety against harmful prompts and avoidance of excessive refusals for benign prompts, measured by the “Not unsafe” and “Not overrefuse” metrics, respectively. The “Not unsafe” metric evaluates the proportion of harmful samples that GPT-4o labels as safe. Conversely, the “Not overrefuse” metric measures the proportion of benign samples that GPT-4o judges as not excessively refused. An example of the evaluation instruction is shown in Figure 6. The overall performance score is computed by averaging the two metrics.
Implementation Details. In the data pre-synthesis stage, the model generates 10 verified safety guidelines. Starting from 10 limited jailbreak examples, we synthesize approximately 5,000 prompts for prompt augmentation to support effective response sampling during fine-tuning. During alignment fine-tuning, responses are sampled based on both the synthesized guidelines and the augmented prompts. For SFT, we select 250 qualified (prompt, response) pairs per data type, yielding 500 samples in total, while for DPO, we select 100 samples per data type, resulting in 200 samples. Only a small subset of synthesized prompts is ultimately used, as many fail to elicit responses meeting the required safety and quality standards.
We adopt LoRA Hu2021LoRALA for efficient fine-tuning. All experiments are conducted for one epoch with a learning rate of $5\times 10^{-4}$, a 10% warm-up ratio, and a cosine scheduler. Inference is performed using vLLM kwon2023efficient with a temperature of 0.6 and a maximum sequence length of 2048. All experiments are run on a single 80GB A800 GPU.